OpenStudy (anonymous):

S3 what means in group

OpenStudy (helder_edwin):

if it is group theory u r talking about. then \(S_3\) is the symmetry group of order 3. That is, \(S_3\) is the group of all permutations defined on the set \(\{1,2,3\}\). It has 3!=6 elements. These 6 elements are: \[ \large \rho_0=\begin{pmatrix} 1 & 2 & 3\\ 1 & 2 & 3 \end{pmatrix} \] \[ \large \rho_1=\begin{pmatrix} 1 & 2 & 3\\ 2 & 3 & 1 \end{pmatrix} \] \[ \large \rho_2=\begin{pmatrix} 1 & 2 & 3\\ 3 & 1 & 2 \end{pmatrix} \] \[ \large \mu_1=\begin{pmatrix} 1 & 2 & 3\\ 1 & 3 & 2 \end{pmatrix} \] \[ \large \mu_2=\begin{pmatrix} 1 & 2 & 3\\ 3 & 2 & 1 \end{pmatrix} \] \[ \large \mu_3=\begin{pmatrix} 1 & 2 & 3\\ 2 & 1 & 3 \end{pmatrix} \]

OpenStudy (helder_edwin):

heres is the answer for one of your last posts @bayanhorani !!

OpenStudy (anonymous):

i have aequation can help me

OpenStudy (anonymous):

about differential equation